OpenStudy (anonymous):

What are the solutions of 3x2 + 4x = -5

OpenStudy (anonymous):

the quantity negative 4 plus or minus i square root of 66 all over 6 the quantity negative 2 plus or minus 2i square root of 66 all over 3 the quantity negative 2 plus or minus i square root of 11 all over 3 the quantity negative 4 plus or minus i square root of 11 all over 6

OpenStudy (anonymous):

for general equation Ax^2 + Bx + C = 0 solution is \[x = \frac{ -B \pm \sqrt{B^{2} - 4*A*C} }{ 2*A }\]

OpenStudy (anonymous):

hope u can use this concept now

OpenStudy (anonymous):

thanks u helped alot